OpenStudy (anonymous):

1 + sec2x sin2x = sec2x Proving trigonometric equations

hartnn (hartnn):

are you sure the question is correct ? its sec4x and not \(\sec^4x\)

hartnn (hartnn):

so, the question really is \(\large 1+\sec^4x\sin^2x=\sec^2x\) ? yes/no ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

col, so lets start by using sec x = 1/cos x plug this into your equation, what u get ?

hartnn (hartnn):

can u ?

OpenStudy (anonymous):

uhm no

hartnn (hartnn):

\(1+\sin^2x/\cos^4x= (\cos^4x+\sin^2x)/\cos^4x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait it's 1+sec^2x not 1+sin^2x

hartnn (hartnn):

oh, that is why i was thinking why i was not able to solve... so, its \(\large 1+\sec^2x\sin^2x=\sec^2x\) ? this is really simple...

OpenStudy (anonymous):

yeah

hartnn (hartnn):

put sec x = 1/cosx and use sin^2x + cos^2x =1

OpenStudy (anonymous):

thanks